OpenStudy (anonymous):

Simplify m^3*n^–6*p^0 use positive exponents

Nnesha (nnesha):

if there is a negative exponent flip it exponent rule \[\huge\rm x^{-m} =\frac{ 1 }{ x^m }\] anything to the 0 power is just equal to one \[\rm (34x)^0 = 1\] \[\rm (Nnesha)^0=1\]